### v1

@@ 3.3 Instrumental polarisation
Regular spectropolarimetric observations started after the installation of the HWP in November 2015. The BST can reduce instrumental polarisation caused by the optical components located between the HWP and the CCD cameras. The residual instrumental polarisation is introduced by the optical elements before the HWP and the combination of the reflective coating of the primary and secondary mirrors. This residual instrumental polarisation was corrected using standard stars with zero degree of polarisation. We correct the Stokes parameters and of the objects for instrumental polarisation using the Stokes parameters and of standard stars with zero degree of polarisation using the following equations:
@@
where and represent both Stokes parameters corrected for the instrumental polarisation.
@@
The 2m mirror’s aluminium reflective coating has to be renewed from time to time. The last two aluminium reflective coatings occurred in 2008 and 2017. Figure 7 shows the Q-U diagram of zero polarisation standard stars in a synthetic V filter before and after recoating of the 2m mirror. A synthetic V filter represents the standard Johnson V-band transmission curve and is used in synthetic photometry to simulate photometric measurements from spectral data. When applied (e.g., via IRAF’s SBAND task), it multiplies the spectrum by the filter transmission curve and integrates the result to estimate the flux or magnitude as if observed through a real V-band filter. This allows a direct comparison between the spectro-polarimetric and polarimetric data.
@@
In Figure 7, different symbols indicate the spectropolarimetric observation modes (with and without slit, before and after the recoating of the 2-m mirror), while the colour represents the modified Julian date (MJD). When a slit is used (triangles and squares), the instrumental polarisation is high and, before the mirror recoating, also highly variable (triangles). In the slitless mode, no significant difference is observed before and after the recoating. The mean Stokes parameters in the synthetic V filter for the slitless mode are and , marked with the black pentagon. The ellipses represent the 1-, 2-, and 3- standard deviations. The diagram also shows a shift of the slitless-mode points (a decrease in Stokes Q), with the latest data lying to the left of the black pentagon, possibly due to changes in instrumental polarisation and the selection of more accurate zero-polarisation standard stars.
@@
Figure 8 shows the polarised spectra of the unpolarised standard star HD 191195 over the wavelength range 4800–8200 Å. No wavelength dependence is observed in the Stokes parameters. The widths of the orange and blue bands indicate the standard deviation from the mean values, calculated over the entire wavelength range. Multiple reflections within the wave plate (e.g., 2003SPIE.4843..437I; 2006AJ....132..433K) can produce the so-called ripple effect; however, no significant (¿0.1%) ripple pattern is detected in the observed Stokes Q and U.
3.3.1 Slit vs. slitless spectropolarimetry
@@
Two types of observations of zero and high polarisation standards were performed - slit and slitless spectropolarimetric observations before and after recoating of the 2m mirror, respectively.
@@
Results of slit- and slitless spectropolarimetric observations before and after recoating of the 2m mirror are presented in Fig. 9 and Fig. 10, respectively. Before the recoating of the 2m mirror, results of slit- and slitless spectropolarimetric observations are practically identical (Fig. 9). After the recoating of the 2m mirror results of slit- and slitless spectropolarimetry observations are different (Fig. 10).
@@
It is known that when the width of a slit is comparable to the wavelength of the incident light, the slit acts as a partial polariser. Even though in our case the slit width is around 150 times the Hα wavelength by using the theory by Slater42, based on microwave wave-guide theory, which also includes the depth of the slit and its material conductivity, we can show that the instrumental polarisation caused only by the slit is in the order of a few percent. This is well explained by AstroSpol-Keller in the book “AstroSpol” ed. by AstroSpol.
@@
In our case, this addition to the instrumental polarisation is most likely due to the combination of the explained slit effect, as well as a small misalignment of the 2m mirror relative to the optical axis, which leads to a slightly deformed star PSF (coma), as well as the narrow metal slit, which cuts the star PSF asymmetrically. This is exactly what you can see on Figures 7 and 10. Before and after the mirror re-aluminization, the Stokes Q has a different value, which is in the order of 1%. Figure 10 even shows how the Stokes U is affected by the instrumental polarisation caused by the coma-slit combination and its wavelength dependence.
Results of slitless spectropolarimetric observations are presented in Section 4. Only when using slitless spectropolarimetric observations we could properly correct for the instrumental polarisation.
@@ 3.4 Correction for the chromatism (wavelength dependence) of the half-wave plate
For correction of the chromatism of the HWP, we used the wavelength dependence function , which is represented in Fig. 11(top panel). This function was derived from the wavelength dependence of the observed position angle of the high polarisation standard star HD 161056. The achromatic function is well described by a sixth-degree polynomial fit:
@@
where and represents Stokes parameter corrected for the chromatism. We used this values of Stokes parameters to calculate . The results of this correction are represented in Fig. 11(bottom panel) where gray data represent the position angle before chromatism correction, while the blue data show it after correction.
3.5 Correction for position angle.
The final correction applied during data reduction is the correction for the position angle. Standard stars with a high degree of polarisation were used to correct the position angle. represents the difference between the observed values in the synthetic V filter and the catalogue value. The correction follows the equation:
where represents the difference between the observed and catalogue values of the position angle. The average value of . The RMS of is , which indicates the instrument’s stability with respect to the position angle correction. In figure 12 we present the observed degree of polarisation in the synthetic V band and . The vertical columns represent catalogue values of the degree of polarisation in the V band, and the width of the column corresponds to the uncertainties in the catalogue values. The horizontal grey line represents the average value of . This offset of P.A. depends on the instrument alignment as well as asymmetric ageing of the mirror coating from season to season. For this reason, it is necessary to determine these P.A. offsets for each observation night or each observation season.
@@
Imaging aperture polarimetry was performed as explained in 2011JQSRT.112.2059B by selecting the aperture at which the reduced Stokes parameters and converge to a well-defined value.
This method differs from the standard aperture photometry with apertures of increasing size in the way we chose the best aperture size. Here, we are looking at which aperture size the ratios between the fluxes in the parallel and the perpendicular beams (hence the measured Stokes parameters) become constant. This aperture is usually smaller than the values where the individual fluxes become asymptotic. Technically, the best aperture is selected using plots like those of Figure 13
The results of the imaging polarimetry measurements through the years of observations with the polarimetric mode of the FoReRo2 instrument are compared with their catalogue values and presented in Figure 14 and Appendix A within Tables 5 and 6 therein.
@@ 4 Spectropolarimetric observations of high polarisation standards
In the spectropolarimetric mode of observation, we observed four unpolarised standard stars: HD 191195, HD 21447, HD 154892, and HD 212311. We also observed ten highly polarised standard stars: HD 7927, HD 25443, HD 161056, HD 19820, HD 204827, BD +25 727, HD 183143, BD +59 389, HD 154445, and HD 155197. The catalogue and observed values of these stars are presented in Table 1 and their polarised spectra are presented in Figure 18. The position angle of the highly polarised standard stars is used as a reference. In Fig. 19, we present 3D maps of these stars. The degree of polarisation is indicated by the length of the orange lines, while the orientation of each line represents the position angle. The colour of each star corresponds to its distance.
Table 1 includes the following information: the observed unpolarised and polarised standard stars, their coordinates, the number of observations for each star, the average value in the synthetic V band, the standard deviation, the degree of polarisation and position angle from catalogues, and references. Two highly polarised standard stars, HD 183143 and HD 204827, exhibited variability during our observations.
| Name | R.A. | Dec. | № of | 1 | 2 | Band | P.D. (%) | P.A. (deg.) | References |
@@
References: (a) 1992AJ....104.1563S; (b) 2020A&A...635A..46P; (c) 1990AJ.....99.1243T;
(d) 1982ApJ...262..732H;
(e) 2021AJ....161...20P; (f) Nordic Optical Telescope3
@@
The average degree of polarisation in the synthetic V band
@@
The standard deviation calculated as in Eq. 8
@@
The average polarisation deviation between the catalogue values and the mean values measured for high-polarisation standard stars in the synthetic V band is (. This value is comparable to the standard deviation for the same set of stars, as listed in Table 1. Therefore, can be interpreted as a representative measure of the instrument’s polarimetric accuracy in the synthetic V band. The stars HD 183143 and HD 204827 were excluded from this calculation due to their known polarimetric variability.
4.1 Searching for variability.
Usually, the standard deviation is used to identify small fluctuations in magnitude and detect flickering (e.g., 2023BlgAJ..38...83Z). In this paper, we apply the same approach to detect and confirm the variability of highly polarised standard stars.
The standard deviation is calculated as:
where is the average degree of polarisation in the synthetic V band, is the number of data points. The calculated in this way includes the variability of the degree of polarisation of the star (if it exists).
@@
Spectropolarimetric observations of HD 183143 with FoReRo2 span a period of two years, from 4 June 2022 to 24 October 2024. The star exhibits an emission line and shows a variable degree of polarisation, indicative of an intrinsic polarisation component. The amplitude of variability in the synthetic V band is . On each of the seven observing nights, another highly polarised standard star was observed alongside HD 183143. The position angle of HD 183143 was corrected using the reference value for the corresponding night. Despite the variation in polarisation degree, the position angle remains stable, with a mean value in the synthetic V band of . Therefore, HD 183143 may still be used as a reference for position angle calibration. It is worth noting that the position angle we derive differs by approximately from the catalogue value, which may indicate long-term variability in the polarisation angle.
1986PASP...98..792D concluded that HD 204827 possesses a component of intrinsic polarisation. Our observations confirm the variability of this star. HD 204827 was observed on two consecutive nights, 2 and 3 May 2025, with position angle calibration performed using a second high-polarisation standard on each night (HD 161056). The applied position angle corrections were and for 2 and 3 May, respectively. HD 204827 exhibits variability in both the degree of polarisation and the position angle. In the synthetic V band, we obtain , on 2 May 2025, and , on 3 May 2025. Despite careful calibration, the derived position angle of HD 204827 differed by nearly between the two nights, with no identifiable instrumental cause. This observed variability suggests that HD 204827 is not a reliable standard for polarimetric calibration, particularly with respect to position angle stability.
@@ 5 FoReRo2 observational polarimetric examples
As illustrative examples, we show a statistical study of and based on the spectropolarimetric observations of polarised stars, spectropolarimetric observations of the recurrent nova RS Oph and the symbiotic star Z And, and the imaging polarimetry of the comet C/2019 Y4 (ATLAS).
5.1 Serkowski law: relationship between and
5.1.1 Fitting the polarisation spectra with the Serkowski law
The Serkowski law is empirical and describes the wavelength dependence of linear polarisation (1975ApJ...196..261S):
where is the peak of the interstellar polarisation at wavelength . The K coefficient describes the shape of the curve, and in Serkowski’s original work (Serkowski et al. 1975), it has a value of 1.15. The parameters of Serkowski’s law fit with catalogue values, and values from this work are presented in Table LABEL:tab.Serkowski. In Fig. 18, polarisation spectra of these high degree of polarisation standards are presented.
@@ 5.1.1 Fitting the polarisation spectra with the Serkowski law
Initially, K = 1.15 was assumed as a standard 1973IAUS...52..145S. An attempt to find a relationship between the parameters of Serkowski law, and can be found in the works of 1982AJ.....87..695W; 1992ApJ...386..562W. In this section, we apply a statistical analysis of and to determine the mean, standard deviation, and data distribution.
1992ApJ...386..562W suggested the relationship:
where is expressed in m. 2017MNRAS.464.4146C, when comparing the K and values presented in his paper with those reported by 1992ApJ...386..562W, reveals a systematic and statistically significant deviation, with the K values reported by 1992ApJ...386..562W, most likely being shifted due to incorrect effective wavelengths for the R and I passbands (sec. 4.2.1. in 2017MNRAS.464.4146C).
Since spectropolarimetric data do not suffer from the problem of properly characterising the photometric system (2017MNRAS.464.4146C), we decided to use only spectropolarimetric data published in the literature, together with data obtained with FR2, to perform a statistical analysis of the values in the plane. We used data published in (2017A&A...608A.146B; 2017MNRAS.464.4146C; 2018A&A...615A..42C; 2022NewA...9701859N; 2023eas..conf.1422N). The coefficients and for nonvariable standard stars are presented in Table LABEL:tab.Serkowski to calculate the mean and standard deviation. For calculations, we used 146 stars, 31 of which were obtained with FR2. The mean of is with a standard deviation of m. The mean of the K coefficient is 1.23 with a standard deviation of 0.32. It is noteworthy that the mean value of K is slightly higher than 1.15, the initial value published by 1973IAUS...52..145S.
In Fig. 15 we present values of Serkowski’s fit of spectropolarimetric data. We present the variable and of RS Oph (with purple) and HD 204827 (with dark blue). With black is presented the sample of Serkowski fit to SNe Ia, taken by (2017ApJ...836...88Z) and Appendix B in (2018A&A...615A..42C). We do not use these values to calculate the mean and standard deviation. In the plane, the spectropolarimetric standard stars (obtained in this work and 2017MNRAS.464.4146C) are located upper to the relationship determined in 1992ApJ...386..562W. Our observations of the standard stars (with red in Figure 15) show the same systematic discrepancy. The orange ellipses present the areas with ; , and . Most of the data is located in . Two of the values for the K coefficient of RS Oph, 2 and 6 days after the most recent outburst, are located outside of . The biggest deviation is observed in the sample of SN Ia (2017ApJ...836...88Z).
@@
Spectropolarimetric observations of the recurrent nova RS Oph following its latest outburst indicate a significant intrinsic component of polarisation (2023A&A...679A.150N). In Table 3, we present the fit of the spectropolarimetric observations of RS Oph using Serkowski’s law. The peak value of the K coefficient is 1.80. From day 2 to day 7 after the 2021 outburst of RS Oph, the K coefficient decreases. It is noteworthy that the peak of the K coefficient within the first week of the most recent outburst of RS Oph corresponds to the newly formed dust during the event. This dust was destroyed within nine days of the 2021 outburst (2023A&A...679A.150N). In Table 3 we present the data with a correlation coefficient higher than 0.75. The variable K coefficient of RS Oph is presented in purple in Figure 15.
The variation in RS Oph is dominated by changes in the Serkowski K coefficient, while remains essentially constant. This suggests that the intrinsic polarisation, due to nova-related dust, affects the shape of the polarisation curve without altering the peak wavelength. In contrast, some supernovae exhibit both high K values and unusually short (e.g., (2017ApJ...836...88Z)), indicating either a different dust grain population or a more complex combination of intrinsic and interstellar polarisation components. Therefore, the behaviour of RS Oph cannot fully account for the polarimetric properties seen in supernovae, but it provides a reference case for transient dust-induced intrinsic polarisation.

### v2

@@ 3.3 Instrumental polarisation
Regular spectropolarimetric observations were initiated after the installation of the HWP in November 2015. The BST can reduce instrumental polarisation caused by the optical components located between the HWP and the CCD cameras. The residual instrumental polarisation is introduced by the optical elements before the HWP and the combination of the reflective coating of the primary and secondary mirrors. This residual instrumental polarisation was corrected using standard stars with zero degree of polarisation. We corrected the Stokes parameters and of the objects for instrumental polarisation using the Stokes parameters and of standard stars with zero degree of polarisation using the following equations:
@@
where and represent both Stokes parameters corrected for instrumental polarisation.
@@
The 2m mirror’s aluminium reflective coating has to be renewed from time to time. The last two aluminium reflective coatings occurred in 2008 and 2017. Figure 7 shows the Q-U diagram of zero polarisation standard stars in a synthetic V filter before and after the re-coating of the 2m mirror. A synthetic V filter represents the standard Johnson V-band transmission curve and is used in synthetic photometry to simulate photometric measurements from spectral data. When applied (e.g., via IRAF’s SBAND task), it multiplies the spectrum by the filter transmission curve and integrates the result to estimate the flux or magnitude as if observed through a real V-band filter. This allows for a direct comparison between the spectro-polarimetric and polarimetric data.
@@
In Figure 7, different symbols indicate the spectropolarimetric observation modes (with and without slit, before and after the re-coating of the 2-m mirror), while the colour represents the modified Julian date (MJD). When a slit is used (triangles and squares), the instrumental polarisation is high and, before the mirror re-coating, also highly variable (triangles). In the slitless mode, no significant difference is observed before and after the re-coating. The mean Stokes parameters in the synthetic V filter for the slitless mode are and , marked with the black pentagon. The ellipses represent the 1-, 2-, and 3- standard deviations. The diagram also shows a shift of the slitless-mode points (a decrease in Stokes Q), with the latest data lying to the left of the black pentagon, possibly due to changes in instrumental polarisation and the selection of more accurate zero-polarisation standard stars.
@@
Figure 8 shows the polarised spectra of the unpolarised standard star HD 191195 over the wavelength range 4800–8200 Å. No wavelength dependence is observed in the Stokes parameters. The widths of the orange and blue bands indicate the standard deviation from the mean values, calculated over the entire wavelength range. Multiple reflections within the wave plate (e.g., Ikeda et al. 2003; Kawabata et al. 2006) can produce the so-called ripple effect; however, no significant (¿0.1%) ripple pattern is detected in the observed Stokes Q and U.
3.4 Slit versus slitless spectropolarimetry
@@
Two types of observations of zero- and high-polarisation standards were performed: slit and slitless spectropolarimetric observations before and after the re-coating of the 2m mirror, respectively. The results of the slit and slitless spectropolarimetric observations before and after the re-coating of the 2m mirror are presented in Figs. 9 and 10, respectively.
@@
Prior to the re-coating of the 2m mirror, the results of slit and slitless spectropolarimetric observations are practically identical (Fig. 9). After the re-coating of the 2m mirror, the results from the slit and slitless spectropolarimetry observations exhibit clear differences (Fig. 10).
@@
It is known that when the width of a slit is comparable to the wavelength of the incident light, the slit acts as a partial polariser. Even though in our case the slit width is around 150 times the Hα wavelength based on the theory by Slater (1942), rooted in microwave wave-guide theory (which also includes the depth of the slit and its material conductivity), we can show that the instrumental polarisation caused only by the slit is in the order of a few per cent. This is aptly explained by Keller (2002)444See Astrophysical Spectropolarimetry, ed. by Trujillo-Bueno et al...
@@
In our case, this addition to the instrumental polarisation is most likely due to the combination of the explained slit effect, as well as a small misalignment of the 2m mirror relative to the optical axis, which leads to a slightly deformed star PSF (coma), as well as the narrow metal slit, which cuts the star PSF asymmetrically. This is exactly what you can see on Figures 7 and 10. Before and after the mirror re-aluminisation, the Stokes Q has a different value, which is in the order of 1%. Figure 10 even shows how the Stokes U is affected by the instrumental polarisation caused by the coma-slit combination and its wavelength dependence. The results of slitless spectropolarimetric observations are presented in Section 4. It is only slitless spectropolarimetric observations that offer the ability to properly correct for the instrumental polarisation.
@@ 3.5 Correction for the chromatism (wavelength dependence) of the half-wave plate
For correction of the chromatism of the HWP, we used the wavelength dependence function , which is represented in Fig. 11 (top panel). This function was derived from the wavelength dependence of the observed PA of the high polarisation standard star HD 161056. The achromatic function is well described by a sixth-degree polynomial fit:
@@
where and represents Stokes parameter corrected for the chromatism. We used these values of Stokes parameters to calculate . The results of this correction are represented in Fig. 11 (bottom panel), where gray data represent the PA before chromatism correction, while the blue data show it after the correction.
3.6 Correcting the position angle
The final correction applied during data reduction is the correction for the PA. Standard stars with a high degree of polarisation were used to correct the PA. Here, represents the difference between the observed values in the synthetic V filter and the catalogue value. The correction follows the equation,
where represents the difference between the observed and catalogue values of the PA. The average value of . The RMS of is , which indicates the instrument’s stability with respect to the PA correction. In Fig. 12 we present the observed degree of polarisation in the synthetic V band and . The vertical columns represent catalogue values of the degree of polarisation in the V band and the width of the column corresponds to the uncertainties in the catalogue values. The horizontal grey line represents the average value of . This offset in the PA depends on the instrument alignment as well as asymmetric ageing of the mirror coating from season to season. For this reason, it is necessary to determine these PA offsets for each observation night or each observation season.
@@ 3.6 Correcting the position angle
Imaging aperture polarimetry was performed as explained in Bagnulo et al. (2011) by selecting the aperture at which the reduced Stokes parameters and converge to a well-defined value. This method differs from the standard aperture photometry with apertures of increasing size in the way we chose the best aperture size. Here, we are looking at which aperture size the ratios between the fluxes in the parallel and the perpendicular beams (and, hence, the measured Stokes parameters) become constant. This aperture is usually smaller than the values where the individual fluxes become asymptotic. Technically, the best aperture can be selected based on the plots in Fig. 13 and similar data. The results of the imaging polarimetry measurements through the years of observations with the polarimetric mode of the FoReRo2 instrument are compared with their catalogue values and presented in Figure 14 and Appendix A within Tables 5 and 6 therein.
@@ 4 Spectropolarimetric observations of high polarisation standards
In the spectropolarimetric mode of observation, we observed four unpolarised standard stars: HD 191195, HD 21447, HD 154892, and HD 212311. We also observed ten highly polarised standard stars: HD 7927, HD 25443, HD 161056, HD 19820, HD 204827, BD +25 727, HD 183143, BD +59 389, HD 154445, and HD 155197. The catalogue and observed values of these stars are presented in Table 1 and their polarised spectra are presented in Figure 18. The PA of the highly polarised standard stars is used as a reference. In Fig. 19, we present 3D maps of these stars. The degree of polarisation is indicated by the length of the orange lines, while the orientation of each line represents the PA. The colour of each star corresponds to its distance.
Table 1 includes the following information: the observed unpolarised and polarised standard stars, their coordinates, the number of observations for each star, the average value in the synthetic V band, the standard deviation, the degree of polarisation and PA from catalogues, and references. Two highly polarised standard stars, HD 183143 and HD 204827, exhibited variability during our observations.
| Name | R.A.1 | Dec.1 | № of | 2 | 3 | Band | P.D. (%) | PA (deg.) | References |
@@
References: (a) Schmidt et al. (1992); (b) Piirola et al. (2020); (c) Turnshek et al. (1990);
(d) Hsu and Breger (1982);
(e) Piirola et al. (2021); (f) Nordic Optical Telescope4
@@
The coordinates were taken from the SIMBAD Astronomical Database.
@@
The average degree of polarisation in the synthetic V band.
@@
The standard deviation calculated as in Eq. 8.
- 4
The average polarisation deviation between the catalogue values and the mean values measured for high-polarisation standard stars in the synthetic V band is (. This value is comparable to the standard deviation
for the same set of stars, as listed in Table 1. Therefore, can be interpreted as a representative measure of the instrument’s polarimetric accuracy in the synthetic V band. The stars HD 183143 and HD 204827 were excluded from this calculation due to their known polarimetric variability.
Usually, the standard deviation is used to identify small fluctuations in magnitude and detect flickering (e.g., Zamanov et al. (2023)). In this work, we applied the same approach to detect and confirm the variability of highly polarised standard stars. The standard deviation is calculated as
where is the average degree of polarisation in the synthetic V band and is the number of data points. The calculated in this way includes the variability of the degree of polarisation of the star (if it exists).
@@
Spectropolarimetric observations of HD 183143 with FoReRo2 span a period of two years, from 4 June 2022 to 24 October 2024. The star exhibits an emission line and shows a variable degree of polarisation, indicative of an intrinsic polarisation component. The amplitude of variability in the synthetic V band is . On each of the seven observing nights, another highly polarised standard star was observed alongside HD 183143. The PA of HD 183143 was corrected using the reference value for the corresponding night. Despite the variation in polarisation degree, the PA remains stable, with a mean value in the synthetic V band of . Therefore, HD 183143 may still be used as a reference for PA calibration. It is worth noting that the PA we derived differs by approximately from the catalogue value, which might indicate long-term variability in the polarisation angle.
Dolan and Tapia (1986) concluded that HD 204827 possesses a component of intrinsic polarisation. Our observations confirm the variability of this star. HD 204827 was observed on two consecutive nights, 2 and 3 May 2025, with PA calibration performed using a second high-polarisation standard on each night (HD 161056). The applied PA corrections were and for 2 and 3 May, respectively. HD 204827 exhibits variability in both the degree of polarisation and the PA. In the synthetic V band, we obtained , on 2 May 2025, and , on 3 May 2025. Despite careful calibration, the derived PA of HD 204827 differed by nearly between the two nights, with no identifiable instrumental cause. This observed variability suggests that HD 204827 is not a reliable standard for polarimetric calibration, particularly with respect to the PA stability.
@@ 5 FoReRo2 observational polarimetric examples
As illustrative examples, we provide a statistical study of and based on the spectropolarimetric observations of polarised stars, spectropolarimetric observations of the recurrent nova RS Oph and the symbiotic star Z And, and the imaging polarimetry of the comet C/2019 Y4 (ATLAS).
5.1 Serkowski’s law: Relation between and
5.1.1 Fitting the polarisation spectra with Serkowski’s law
Serkowski’s law is empirical and describes the wavelength dependence of linear polarisation (Serkowski et al., 1975), following
where is the peak of the interstellar polarisation at wavelength . The K coefficient describes the shape of the curve and based on Serkowski’s original work (Serkowski et al. 1975), it has a value of 1.15. The parameters of the Serkowski law fit are presented in Table LABEL:tab.Serkowski. In Fig. 18, the polarisation spectra of these high degree of polarisation standards are presented.
@@ 5.1.1 Fitting the polarisation spectra with Serkowski’s law
Initially, K = 1.15 was assumed as a standard in the study by Serkowski (1973). An attempt to find a relation between the parameters of Serkowski’s law, , and can be found in the works of Wilking et al. (1982); Whittet et al. (1992). In this section, we applied a statistical analysis of and to determine the mean, standard deviation, and data distribution. Whittet et al. (1992) suggested the following relation,
where is expressed in m. Cikota et al. (2017) compared the K and values presented in his paper with those reported by Whittet et al. (1992), revealing a systematic and statistically significant deviation, with the K values reported by Whittet et al. (1992), most likely being shifted due to incorrect effective wavelengths for the R and I passbands (see Section 4.2.1. in Cikota et al. (2017)).
Since spectropolarimetric data do not suffer from the problem of properly characterising the photometric system (Cikota et al., 2017), we decided to use only spectropolarimetric data published in the literature, together with data obtained with FR2, to perform a statistical analysis of the values in the plane. We used data published in (Bagnulo et al., 2017; Cikota et al., 2017, 2018; Nikolov, 2022; Nikolov and Kaltcheva, 2023). The coefficients and for non-variable standard stars are presented in Table LABEL:tab.Serkowski. To calculate the mean and standard deviation, we used 146 stars, 31 of which were obtained with FR2. The mean of is with a standard deviation of m. The mean of the K coefficient is 1.23 with a standard deviation of 0.32. It is noteworthy that the mean value of K is slightly higher than 1.15, the initial value published by Serkowski (1973).
In Fig. 15, we present values of Serkowski’s fit of the spectropolarimetric data. We present the variable and of RS Oph (with purple) and HD 204827 (with dark blue). The sample of SNe Ia (Zelaya et al., 2017) is shown in black, with details given in Appendix B of Cikota et al. (2018). We did not use these values to calculate the mean and standard deviation. In the plane, the spectropolarimetric standard stars (obtained in this work and Cikota et al. (2017)) are located above the relation determined in Whittet et al. (1992). Our observations of the standard stars (with red in Figure 15) show the same systematic discrepancy. The orange ellipses present the areas with ; , and . Most of the data is located in . Two of the values for the K coefficient of RS Oph (2 and 6 days after the most recent outburst) are located beyond . The biggest deviation is observed in the sample of SN Ia (Zelaya et al., 2017).
@@
Spectropolarimetric observations of the recurrent nova RS Oph following its latest outburst indicate a significant intrinsic component of polarisation (Nikolov et al., 2023). In Table 3, we present the fit of the spectropolarimetric observations of RS Oph using Serkowski’s law. The peak value of the K coefficient is 1.80. From day 2 to day 7 after the 2021 outburst of RS Oph, the K coefficient decreases. It is noteworthy that the peak of the K coefficient within the first week of the most recent outburst of RS Oph corresponds to the newly formed dust during the event. This dust was destroyed within nine days of the 2021 outburst (Nikolov et al., 2023). In Table 3 we present the data with a correlation coefficient higher than 0.75. The variable K coefficient of RS Oph is presented in purple in Figure 15.
The variation in RS Oph is dominated by changes in the Serkowski K coefficient, while remains essentially constant. This suggests that the intrinsic polarisation, due to nova-related dust, affects the shape of the polarisation curve without altering the peak wavelength. In contrast, some supernovae exhibit both high K values and unusually short (e.g., (Zelaya et al., 2017), indicating either a different dust grain population or a more complex combination of intrinsic and interstellar polarisation components. Therefore, the behaviour of RS Oph cannot fully account for the polarimetric properties seen in supernovae, but it does provide a reference case for transient dust-induced intrinsic polarisation.